harleigh:

Cheyanne plans to put $150 into a savings account. She can place her money into an account represented by f(x) = 5x + 150, or into another account represented by g(x) = 150(1.03)x. Which account has the highest value in 3 years? Which account has the highest value in 10 years?

jhonyy9:

hi ! any idea ?

jhonyy9:

there are f(x) and g(x) ,yes ? how you think what are the values of x ?

jhonyy9:

are you here please ? so please collaborate

dude:

Since you are gone You would have to substitute 3 into both equations to see which is greater at both. Then substitute 10 into both to see which is greater at 10 years For 3 years: \(f(x) = 5(3) + 150=> \boxed{165}\) \(g(x) = 150(1.03)^{3}\approx> \boxed{163.909}\) f(x) is greater here (After 3 years) For 10 years: \(f(x) = 5(10) + 150=> \boxed{200}\) \(g(x) = 150(1.03)^{10}\approx> \boxed{201.587}\) g(x) is greater (After 10 years)